Applying Cauchy’s mean value theorem to the function f and g defined as f(x)=x2 and g(x)=x for all x∈[a,b], gives

1
Expert's answer
2022-03-14T18:50:59-0400

Solution


Cauchy's Mean Value Theorem is defined as:


"\frac{{f'\left( c \right)}}{{g'\left( c \right)}} = \frac{{f\left( b \right) - f\left( a \right)}}{{g\left( b \right) - g\left( a \right)}}"


Here, given that


"f(x)=x^2" and "g(x)=x" and the interval is "[a, b]"


Then,


"f'(x)=2x\\
f'(c)=2c\\" and "g'(x)=1\\
g'(c)=1\\"


Therefore, as per Cauchy's Mean Value Theorem


"\frac{{f'\left( c \right)}}{{g'\left( c \right)}} = \frac{{f\left( b \right) - f\left( a \right)}}{{g\left( b \right) - g\left( a \right)}}"


"\frac{2c}{1} = \frac{b^2-a^2}{b-a}"


"\frac{2c}{1} = \frac{(b-a)(b+a)}{b-a}"


"2c=a+b"


Hence the Cauchy's Mean Value Theorem gives,


"a+b=2c"